OpenStudy (anonymous):

Solve by taking square roots. (y + 5)2 − 50 = 0 Smaller number is _____________ Larger number is _______________

OpenStudy (anonymous):

Assuming the problem is: (y+5)^2 - 50 = 0; You get: (y+5)^2 = 50 (y+5) = sqrt(50) y+5 = +/- 7.07 y = 2.07 or -12.07

OpenStudy (mertsj):